OpenStudy (anonymous):

Find three cube roots of 64i

OpenStudy (kc_kennylau):

Express this in the form of a(cosx+isinx) first

OpenStudy (anonymous):

4(cos+isinx) ?

OpenStudy (kc_kennylau):

I mean express 64i in the form of a(cosx+isinx)

OpenStudy (anonymous):

\[ \sqrt[3]{64 i}=\sqrt[3]{64 e^{\frac{i \pi }{2}}}=\\4 e^{\frac{i \pi }{6}}=4 \left(\cos \left(\frac{\pi }{6}\right)+i \sin \left(\frac{\pi }{6}\right)\right)=\\4 \left(\frac{\sqrt{3}}{2}+\frac{i}{2}\right)=2 \sqrt{3}+2 i \]